# Subwavelength arrays of quantum emitters: nonlinearities enter the weak-drive regime, and lead to correlated subradiant states

### v3

@@ Abstract
Quantum emitter arrays have emerged as important platforms in which strong light-matter interactions can be achieved and precisely controlled. In subwavelength regimes, they are characterised by a manifold of subradiant eigenstates, which can host novel quantum many-body states and are promising for several applications. Nevertheless, their excitation under a weak far-field driving is prevented by their subradiant nature. Consequently, the weak-drive regime has been widely regarded as a linear regime, well described by classical equations. In this work, we find that, suprisingly, the weak-drive regime is instead strongly non-linear for regular geometries, due to the important contribution of higher-order, resonant and momentum-conserving processes exciting subradiant states. Using a Dynamical Mean-Field Theory (DMFT) we show that these processes lead to a steady state of interacting subradiant excitation pairs, with quantum correlations of two-mode squeezing and long-range character, realising a novel strongly-correlated driven-dissipative state. These findings have potential impact for non-linear optics and quantum metrology.
@@
Understanding the collective behaviour of an ensemble of quantum emitters coupled to the electromagnetic radiation is a long-standing problem in atomic physics and quantum optics [1, 2, 3, 4]. When the inter-emitter distance is made comparable or smaller than the wavelength of the radiative transition, their dipolar interactions become strong and the emission becomes collective, resulting in a strongly-interacting and dissipative quantum many-body problem. These subwavelength conditions can lead to spectacular collective effects such as cooperative Lamb shifts [5, 6], superradiance and subradiance [1, 4].
@@
The problem has recently received renewed attention as different experimental platforms – such as ultracold atoms [7, 8] and semiconductor dipolar excitons [9, 10, 11] – are now entering this strongly-interacting regime. Combined with their ability to arrange the emitters in precise geometries, this results in an exquisite control of the light-matter interaction and of the emerging collective behaviour. These capabilities have recently led to remarkable applications, such as the realisation of single-layer mirrors of atoms [12, 13, 8] – where coherent and spatial control of their reflectivity was also achieved [7] – as well as of excitons [14, 11, 15, 10], contributing towards the realisation of metasurfaces that can control the quantum properties of light [16, 7, 17].
@@
Nevertheless, applications, as well as most theoretical understanding, have been so far limited to the regime of linear response. Harnessing the non-linearities of quantum emitters arrays offers the potential for far-reaching consequences for non-linear quantum optics [18, 19, 17, 20], quantum metrology [21], as well as for quantum computing, where atom arrays have emerged as promising architectures [22, 23, 24]. In particular, these platforms could host novel quantum many-body states due to their unique interplay of dissipation and interactions, which could be harnessed as resources in those fields.
The subradiant eigenstates of these systems [25, 26, 27] provide an interesting basis with which to realise correlated and/or entangled states: their coupling to free photons is strongly suppressed resulting in a largely suppressed radiative decay. While their single-excitation regime is well understood [25], and has led to several proposals for applications for photon storage [28, 29, 25, 30, 31, 32, 33, 34] and quantum metrology [35, 36, 37], only little progress has been made on understanding their many-body physics [25, 21, 27].
On the other hand, due to their subradiant nature, these states are also difficult to excite and manipulate [25, 38, 39]. Under a weak far-field driving, with which coherent control could be achieved and for which heating effects could be contained, the contribution of subradiant states is widely expected to be negligible. This expectation had as a consequence that the weak-drive regime has been long regarded as a linear regime accurately described by classical equations of motion [40, 41, 42, 43, 44, 12, 45, 46, 47, 48, 49], which would become an exact description in the limit of weak drive intensity [40, 41, 42, 43]. Alternative proposals for subradiant-states manipulation were made [39, 50, 29, 32, 33, 34], but these require additional, often near-field, geometrical control, and are limited to the single-excitation regime.
In this work we show that, by contrast with the above established expectations, the weak-drive regime of a large array of quantum emitters is instead surprisingly strongly non-linear, owing to the important effect of resonant, momentum-conserving processes, exciting pairs of single-particle subradiant modes. These processes have been overlooked in previous analyses [40, 41, 42, 43, 44, 12, 45, 46, 47, 48, 49]. We show that they lead to a steady-state of interacting subradiant excitation pairs.
We first demonstrate the importance of non-linearities beyond classical theories by deriving a weakly non-linear theory of the steady-state, in which these are accounted for perturbatively in the drive strength. We show that such a perturbative approach leads to a divergence of the steady state for a drive strength that vanishes in the limit of a large number of emitters. As a result, for large arrays the weak-drive steady-state is non-perturbative with respect to the prediction of classical theories. These approaches [40, 41, 42, 43, 44, 12, 45, 46, 47, 48, 49], as well as perturbative ones, are therefore inadequate.
@@
Then, we study the resulting steady-state under continuous driving using the non-perturbative Dynamical Mean-Field Theory (DMFT) for Markovian spin systems recently introduced in [51, 52]. This has been previously applied only to the regime of strong drive intensities [51], in which the steady-state is a trivial infinite temperature state — albeit with an interesting fluorescence spectrum [51].
@@
Using DMFT, we predict a steady-state of interacting subradiant excitation pairs. We show that this is characterised by quantum correlations with two-mode squeezing and long-range character, and that these survive heating effects due to driving, dissipation and interactions. Moreover, we demonstrate that the effects of non-linearities in the weak drive regime are strong: namely, we show that the non-linear steady-state population is comparable in magnitude with the linear component predicted by classical theories.
Our findings can be realised in current experiments with dipolar excitons [9], and could be realised in upcoming ones with ultracold atoms [8, 7] that will soon operate in suitable regimes. They open new exciting avenues for applications: The non-linear regime at weak driving uncovered can be harnessed for non-linear optics [18, 19, 17], for example for the generation of entangled photon pairs. Moreover, the predicted steady-state has potential for quantum metrology, owing to its multi-mode squeezing correlations which are sought after in this field [53].
@@
Figure 1: (a) A schematic of a 1D array of two-level emitters, with transition wavevector $k_{0}$ and lattice spacing $a$, under external drive. The driving electric field vector E is oriented along the chain, inducing the orientation of the atomic dipoles indicated by an arrow. The emitters are subject to coherent dipole-dipole interactions $V_{ij}$ and collective decay processes with rates $\Gamma_{ij}$, both of which are long-ranged. (b) The dispersion relation $V_{k}$ (solid line) and decay rates $\Gamma_{k}$ (dash-dot line) of collective single-particle modes for $N\rightarrow\infty$ emitters, as functions of momentum $k$, for the subwavelength lattice spacing $k_{0}a=2$: the modes for $\left\lvert k\right\rvert>k_{0}$ in the non-grayed area correspond to subradiant modes with zero decay rates, that are linearly decoupled from the far-field drives; the illustration instead depicts a non-linear process, in which two drive photons effectively scatter resonantly into a pair of subradiant modes. (c) The number of steady state excitations $n_{c}(k)=\langle{\sigma_{k}^{+}\sigma_{k}^{-}}\rangle_{c}$ calculated in DMFT/NCA for $N\rightarrow\infty$, as a function of momentum and drive detuning $\Delta$ (modes dispersion $V_{k}$ is superimposed in yellow). Here the drive strength is $\Omega/\Gamma=0.8$, the lattice spacing the same as in (b), and the subscript “c” indicates a connected average. In particular, it shows that the modes in the non-grayed region are resonantly populated.
@@
Model. We consider $N$ two-level emitters ordered in the one-dimensional (1D) geometry illustrated in Fig. 1 (a). We focus on this set up for simplicity, but most of our conclusions and methods can also be extended to more general geometries, and to some extent to two dimensions. The emitters are illuminated with a uniform plane wave, and are coupled to the free-space electromagnetic vacuum, giving rise both to coherent dipole-dipole interactions and collective decay.
Such a system can be described by the Markovian master equation [2]
@@
Here $\sigma_{l}^{\alpha}$ with $\alpha=x,y$ or $z$ are the Pauli matrices on site $l$, and ${\sigma}_{l}^{\pm}={\sigma}_{l}^{x}\pm{\rm i}{\sigma}_{l}^{y}$ the raising and lowering operators. $\Delta=\omega_{d}-\omega_{0}$ is the detuning of the driving field frequency from the two-level transition energy $\omega_{0}$, and $\Omega=2\mathbf{d}\cdot\mathbf{E}/\hbar$ is the Rabi coupling, given by the vector of transition dipole moments $\mathbf{d}$ and by the driving electric-field vector $\mathbf{E}$. Such field is assumed uniform and oriented along the array, leading to the rates of coherent dipolar interactions and collective decay
@@
A key quantity for these couplings is the ratio between the lattice spacing $a$ and the emitter transition wavelength $\lambda_{0}$, entering via the parameter $2\pi a/{\lambda_{0}}=k_{0}a$, where $k_{0}$ is the transition wavevector. This controls the relative amplitude and range of the couplings, such that for decreasing $k_{0}a$ they become longer-ranged and the coherent interactions increase in magnitude, while for increasing values the regime of independent emitters is approached. Here we consider intermediate values, for which both couplings are of the same order, and thus interesting driven-dissipative dynamics is expected. Hereafter we assume units of $\Gamma=1$ and $\hbar=1$.
@@
In the particular case in which a single excitation is prepared in the system, and in absence of driving $\Omega=0$, such equations become equivalent to a problem of non-interacting bosons. This is obtained by replacing the spin operators with bosonic ones $\sigma^{-}_{j}\rightarrow a_{j}$ [54, 25, 55], via a Holstein-Primakoff representation of the spin. Such a problem can be solved exactly in terms of eigenmodes of collective excitations, with frequencies shifted from those of individual emitters of $V_{\alpha}$, and with decay rates $\Gamma_{\alpha}$, where $\alpha$ is the mode index. For a non-zero but weak driving, one may also approximate (1)-(3) with such single-particle non-interacting bosonic theory upon adding a linear drive term, provided that the individual population of the collective modes remains weak. For a resonant monocromatic drive, this typically holds for a drive strength smaller than the driven mode decay rate $\Omega\ll\Gamma_{\alpha}$.
@@
Additionally, in the subwavelength regime in which the emitters spacing is smaller than their transition wavelength, modes with decay rates much smaller than independent emitters $\Gamma_{\alpha}\ll\Gamma$ appear, that are called subradiant [4, 25]. These arise due to destructive interference of the field emitted by different emitters. Despite the long lifetimes, by the same interference mechanism subradiant modes are also linearly decoupled from far-field drives, thus they are not easily excited under a weak driving [25, 38, 39]. We also remark that this discussion is not specific to the case of arrays, but also extends to disordered ensembles as realised in cold-atoms experiments [56, 57, 31, 58, 49]: these are also described by Eqs. (1)-(3) upon specifying the appropriate couplings $V_{ij}$ and $\Gamma_{ij}$.
@@
For the above reasons, even in the subwavelength regime and in the presence of subradiant states, the weak-drive regime has been widely regarded as a linear regime that is accurately described by classical equations of motion [40, 41, 42, 43, 44, 12, 45, 46, 47, 48, 49], equivalent to the non-interacting bosons description. These have been expected to become exact descriptions in the limit of vanishing drive intensity [40, 41, 42, 43], without any distinction being made between the cases of disordered ensembles and regular arrays. By contrast, we will show that, in the case of large subwavelength emitters arrays, the weak-drive regime is strongly non-linear due to the important contribution of higher-order resonant and momentum conserving processes, which has been overlooked in previous analyses [40, 41, 42, 43, 44, 12, 45, 46, 47, 48, 49].
@@
In the case of arrays, lattice momentum k is a good quantum number for $N\rightarrow\infty$, and the bosonic eigenmodes correspond to reciprocal-space modes $a_{\textbf{k}}=\sum_{j}{\rm e}^{{\rm i}\textbf{k}\cdot\textbf{ r}_{j}}a_{j}/\sqrt{N}$, with dispersion relation and decay rates given by the Fourier transforms of (4) and (5) $V_{\textbf{k}}$ and $\Gamma_{\textbf{k}}$, shown in Fig. 1 (b) for the considered geometry. In the subwavelength regime, an entire manifold of subradiant eigenmodes arises, which have decay rates strongly suppressed by the number of emitters. These decay rates have been found to depend polynomially on $N$ for open boundary systems $\Gamma_{\textbf{k}}\sim N^{-\alpha}$, with $\alpha$ seemingly universal for the most subradiant modes with $\alpha=3$ in 1D, and $\alpha=6$ for a 2D square lattice [25]. It was also shown that the residual decay is mainly due to scattering off the array boundaries, and that for periodic boundary conditions in 1D the suppression becomes exponential $\Gamma_{\textbf{k}}\sim\exp(-N)$.
@@
This suppression can be understood by such modes having an energy much smaller than free-space photons with the same longitudinal momentum – the momentum along the array –, thus their excitations cannot decay into single photons conserving energy and momentum [25, 38]. From this argument, it follows that this happens whenever the mode dispersion intersects the light-cone within the Brillouin zone [25], corresponding to the condition $k_{0}a\leq\pi$ in 1D and $k_{0}a\leq\sqrt{2}\pi$ for a 2D square lattice, and that those modes appear at large momenta $\left\lvert\textbf{k}\right\rvert>k_{0}$, as in the case shown in Fig. 1 (b).
@@
Failure of perturbative approaches. While the direct coupling of a far-field drive with subradiant modes is suppressed, resonant excitation of these modes can still happen through higher-order processes in perturbation theory, mediated by the non-linearities of the atoms.
@@
These processes can be identified by expanding a Holstein-Primakoff representation of the spin in terms of bosons, under the assumption that the drive is sufficiently weak such that the steady state has a low density of excitations around the state with all spins pointing down — more details in [59]. One can then expand simultaneously in the drive and in the non-linear terms in a time dependent perturbation theory, and retain only resonant processes that are not energetically suppressed. For the geometry illustrated in Fig. 1 (a), these involve a pair of subradiant modes with $V_{\pm k}=\Delta$. The lowest-order such processes, in momentum space, read $(\Omega/\sqrt{N}a^{\dagger}_{-k}a^{\dagger}_{k}a_{k=0})_{t_{2}}({\sqrt{N}\Omega}a^{\dagger}_{k=0})_{t_{1}}$ and $({V_{k}}{/N}a^{\dagger}_{-k}a^{\dagger}_{k}a_{k=0})_{t_{3}}({\sqrt{N}\Omega}a^{\dagger}_{k=0})_{t_{2}}({\sqrt{N}\Omega}a^{\dagger}_{k=0})_{t_{1}}$, where $a$ is the boson annihilation operator. These excite a pair of subradiant modes with momenta (k,k)(-k,k), via an intermediate (virtual) excitation of the $k=0$ mode. Effectively, they convert two drive photons – the $\Omega^{2}$ factors – into a pair of subradiant excitation, conserving energy and momentum. This is illustrated in panel (b) of Fig. 1. We note that, as opposed to the case of disordered ensembles, the periodic geometry of arrays also allows the matrix element of these processes – between the vacuum and the subradiant eigenstates – to be significant.
@@
To estimate the effect of these processes, we derived [59] a weakly-non-linear theory of the steady-state, where they are accounted for perturbatively in the drive strength with respect to the non-interacting steady-state. Neglecting the fluctuations of the $k=0$ mode — which is directly driven — around its average value leads to an effective model of a generalised parametric amplifier for the subradiant-modes pair, with Hamiltonian $H_{k}=\delta_{-k}(a_{k}^{\dagger}a_{k}+a_{-k}^{\dagger}a_{-k})+\tilde{\delta}_{-k}(a_{k}^{\dagger}a_{-k}+a_{-k}^{\dagger}a_{k})+\lambda_{-k}(a_{k}^{\dagger}a_{k}^{\dagger}+a_{-k}^{\dagger}a_{-k}^{\dagger}+\text{h.c.})+\tilde{\lambda}_{-k}(a_{k}^{\dagger}a_{-k}^{\dagger}+\text{h.c.})$, where all coefficients are independent of $N$ and of leading order $\Omega^{2}$ in the drive strength. The dissipator at leading order is simply obtained by replacing $\sigma^{-}\rightarrow a$ in (3) and is diagonal in momentum, where further corrections are negligible since the decay rate of subradiant modes is already suppressed for large $N$.
The result of this analysis is that such a perturbative approach can only describe the steady state at incredibly weak drive strengths. In fact, for $\Omega\gtrsim\Gamma_{\left\lvert k\right\rvert}^{1/2}$ the steady state develops a well-known parametric instability [60, 59], accompanied by a divergence of its populations. For the most subradiant modes, this corresponds to $\Omega\gtrsim N^{-\alpha/2}$ for open boundary conditions, and $\Omega\gtrsim\exp(-N/2)$ for periodic ones. Beyond this threshold, the effect of non-linearities beyond non-interacting (classical) theories is therefore non-perturbative, and these theories as well as perturbative approaches are inadequate. We note that such a threshold can be very small already for moderate $N$, and that for $N\rightarrow\infty$ the weak drive steady-state $\Omega\rightarrow 0$ is genuinely non-perturbative with respect to the undriven steady-state at $\Omega=0$ (the vacuum).
@@
A correlated steady-state of subradiant excitation pairs. To capture the steady state in the non-linear (many-body) weak-drive regime discussed, we used the non-perturbative Dynamical Mean-Field Theory (DMFT) for Markovian spin systems developed in [51, 52], for homogeneous steady states and in the thermodynamic limit $N\rightarrow\infty$. This is a bosonic version of DMFT, that maps the lattice model (1) onto an effective model of a single-site coupled to an effective classical field and to a non-Markovian environment, which are self-consistently determined, where the former is the analogous of a Gutzwiller mean-field theory and the latter captures correlations beyond such theory. The effective model is reported in the Supplemental Material [59]. There, we also show that DMFT becomes exact in the non-interacting limit — in addition to the limit of independent emitters — which is the relevant reference state in the weak-drive regime.

### v4

# How Subradiance Enables Nonlinearity in Weakly Driven Quantum Arrays

@@ Abstract
Harnessing the nonlinear response of a medium is essential for applications including frequency conversion and light amplification, as well as for the generation of quantum many-body correlations of light or matter. However, achieving these effects typically requires high drive intensities and thick samples, which induce undesired heating effects that typically suppress quantum correlations. In this work, we demonstrate that atom-thin arrays of quantum emitters exhibit a robust nonlinear response even at arbitrarily weak drive intensities. This discovery challenges the long-held assumption that weakly driven ensembles behave classically; instead, we reveal that subradiant states provide a dominant nonlinear contribution that persists in the low-intensity limit. Using a Dynamical Mean-Field Theory (DMFT) approach, we predict that these nonlinearities generate a quantum-correlated steady state composed of interacting pairs of subradiant excitations, characterized by long-range correlations and multi-mode squeezing. Our findings establish a new frontier for nonlinear quantum optics at minimal power, and provide a scalable protocol for preparing multimode squeezing, offering potential for applications in quantum metrology.
@@
Understanding the collective behaviour of an ensemble of quantum emitters coupled to the electromagnetic radiation is a long-standing problem in atomic physics and quantum optics [16, 37, 1, 25]. When the inter-emitter distance is made comparable or smaller than the wavelength of the radiative transition, their dipolar interactions become strong and the emission collective, resulting in a strongly-interacting and dissipative quantum many-body problem. These subwavelength conditions lead to spectacular collective effects such as cooperative Lamb shifts [20, 31], superradiance and subradiance [16, 25].
@@
The problem is receiving renewed attention, as different experimental platforms – ultracold atoms [62, 52] and semiconductor dipolar excitons [35, 59, 5] – are now entering the strongly-interacting regime. Combined with the experimental ability to arrange quantum emitters in precise geometries, the achievement of strong coupling results in an exquisite control of the light-matter interaction and of the emerging collective behaviour. These capabilities have recently led to remarkable applications, such as the realisation of single-layer mirrors of atoms [9, 60, 52] – with coherent and spatial control of their reflectivity [62] – as well as of excitons [68, 5, 70, 59], contributing towards the realisation of metasurfaces that can control the quantum properties of light [7, 62, 61].
@@
Applications, as well as most theoretical understanding, have been so far limited to the regime of linear response. Harnessing the nonlinear properties of quantum emitters arrays offers the potential for far-reaching consequences for quantum nonlinear optics [33, 39, 61, 65], as well as for quantum metrology [28], and quantum computing (where atom arrays are one of the most promising architectures [41, 11, 10]).
The subradiant eigenstates [4, 69, 50] of these systems – interesting as quantum memories [46, 17, 4, 43, 19, 6, 51, 13] and for quantum metrology [44, 18, 48] – provide an attractive basis where to probe strong nonlinear effects [4, 28, 50], owing to their heavily suppressed decay rates. Nevertheless, these are difficult to excite and manipulate. In particular, they cannot be excited with a weak, (linearly-) resonant electromagnetic driving field [4, 27, 67] (they are dark states). This fact prevents the possibility of driving them weakly into nonlinear regimes, without generating substantial heating that typically largely suppresses quantum correlations.
The view that subradiant state are negligible has had the far-reaching consequence that the weak-drive regime of emitter ensembles has long been regarded as one of linear response, accurately described by classical equations of motion [53, 29, 36, 66, 8, 9, 45, 63, 49, 24, 26], which have been viewed as an exact description of the limit of weak drive intensity [53, 29, 36, 66].
In this work, we show that, by contrast, the weak-drive regime of a large array of quantum emitters is non-perturbative and surprisingly nonlinear. This is due to resonant “parametric driving” processes, involving pairs of single-particle subradiant modes. We first illustrate their effect by treating them perturbatively with respect to noninteracting theories [53, 29, 36, 66, 8, 9, 45, 63, 49, 24, 26]. We show that this amounts to constructing a perturbation theory around a model of a parametric amplifier [58], which becomes dynamically unstable beyond a threshold driving strength thereby signalling a non-perturbative steady state. Because this threshold decreases to zero with increasing emitter number, the established classical theories  [53, 29, 36, 66, 8, 9, 45, 63, 49, 24, 26] prove inadequate to describe the weak-drive behaviour of large arrays.
@@
Then, we study the resulting steady-state under continuous driving, using the non-perturbative Dynamical Mean-Field Theory (DMFT) for Markovian spin systems recently introduced [55, 54]. This has been so far applied only to the regime of strong drive intensities [55] – in which the steady-state is a trivial infinite temperature state, albeit with an interesting fluorescence spectrum [55]: here we address the more challenging  [42, 49] weak-driving regime, which required several technical improvements (discussed later in the main text, as well as in SM Notes C and E).
@@
Using DMFT, we find that the weak-drive nonlinearities lead to a steady state of interacting pairs of subradiant excitations. This is characterised by quantum correlations with multimode squeezing and long-range character, which are robust to the heating caused by driving, dissipation and interactions. Moreover, we demonstrate that these nonlinear effects are strong: the nonlinear steady-state population is comparable with the linear component predicted classically. This confirms the inadequacy of classical theories [53, 29, 36, 66, 8, 9, 45, 63, 49, 24, 26], and indicates a high conversion efficiency of drive photons into the nonlinearly-populated modes, relevant for nonlinear optics.
Our findings can be realised in current experiments with dipolar excitons [35], and in upcoming ones with ultracold atoms [52, 62], which will soon operate in suitable regimes. They open a new frontier for nonlinear quantum optics at minimal power and with atom-thin media, avoiding the significant heating generated in typical strong-driving regimes, thereby allowing a much larger degree of quantum correlations and quantum non-linear effects; for example, we discuss how our setup could enable enhanced generation of entangled photons. Our work also provides a simple scheme to generate multimode squeezing correlations, offering potential for applications in quantum metrology, where such schemes and correlations are highly sought after [40].
@@
Figure 1: (a) A schematic of a 1D array of two-level emitters, with transition wavevector $k_{0}$ and lattice spacing $a$, under external driving. The driving electric field vector E is oriented along the chain, inducing the orientation of the atomic dipoles indicated by an arrow. The emitters are subject to coherent dipole-dipole interactions $V_{ij}$ and collective decay processes with rates $\Gamma_{ij}$, both of which are long-ranged. (b) The dispersion relation $V_{k}$ (solid line) and decay rates $\Gamma_{k}$ (dash-dot line) of collective single-particle modes for $N\rightarrow\infty$ emitters, as functions of momentum $k$, for the subwavelength lattice spacing $k_{0}a=2$: the modes for $\left\lvert k\right\rvert>k_{0}$ – in the non-grayed area – correspond to subradiant modes with zero decay rates (in the thermodynamic limit), that are linearly decoupled from the far-field; the illustration instead depicts a nonlinear process, in which two drive photons scatter resonantly into a pair of subradiant modes. (c) The number of steady state excitations $n_{c}(k)=\langle{\sigma_{k}^{+}\sigma_{k}^{-}}\rangle_{c}$ calculated in DMFT/NCA for $N\rightarrow\infty$, as a function of momentum and drive detuning $\Delta$ (the dispersion relations is superimposed in yellow). Here the drive strength is $\Omega/\Gamma=0.8$, the lattice spacing is the same as in (b). It shows that the subradiant modes are resonantly populated.
@@
Model. We consider $N$ two-level emitters ordered in a one-dimensional (1D) geometry, and specifically the one illustrated in Fig. 1 (a). Most of our conclusions and methods can nevertheless be extended to more general 1D geometries, as well as to two dimensions. The emitters are illuminated with a uniform plane wave, and are coupled to the free-space electromagnetic vacuum, giving rise both to coherent dipole-dipole interactions and to collective decay.
Such system can be described by the Markovian master equation [37]
@@
Here $\sigma_{l}^{\alpha}$ with $\alpha=x,y$ or $z$ are the Pauli matrices on site $l$, and ${\sigma}_{l}^{\pm}={\sigma}_{l}^{x}\pm{\rm i}{\sigma}_{l}^{y}$ the raising and lowering operators. $\Delta=\omega_{d}-\omega_{0}$ is the detuning of the driving-field frequency from the two-level transition energy $\omega_{0}$, and $\Omega=2\mathbf{d}\cdot\mathbf{E}/\hbar$ is the Rabi frequency – indicating the strength of the driving –, given by the vector of transition dipole moments $\mathbf{d}$ and by the driving electric-field vector $\mathbf{E}$. Such field is assumed uniform and oriented along the array (see Fig. 1 (a)), leading to the rates of coherent dipolar interactions and collective decay
@@
A key parameter determining these couplings is $k_{0}a=2\pi a/{\lambda_{0}}$, namely the ratio between the lattice spacing $a$ and the emitter transition wavelength $\lambda_{0}$ – where $k_{0}$ is the transition wavevector. It controls the relative amplitude and range of the couplings, such that for decreasing $k_{0}a$ they become longer-ranged and the coherent interactions increase in magnitude, while for increasing values the regime of independent emitters is approached. Here we consider intermediate values, for which both couplings are of the same order, and thus interesting driven-dissipative dynamics is expected. Hereafter, we assume units of $\Gamma=1$ and $\hbar=1$.
@@
In the particular case in which a single excitation is prepared in the system, and in absence of driving $\Omega=0$, such equations become equivalent to a problem of noninteracting bosons. This is obtained by replacing the spin operators with bosonic ones $\sigma^{-}_{j}\rightarrow a_{j}$ [47, 4, 64], via a Holstein-Primakoff representation of the spin. It can be solved exactly in terms of eigenmodes of collective excitations, with frequencies shifted from those of individual emitters of $V_{\alpha}$, and with decay rates $\Gamma_{\alpha}$, where $\alpha$ is the mode index. For a non-zero yet weak driving, one may also approximate (1)-(3) with such single-particle bosons theory – upon adding a linear drive term –, provided that the individual population of the collective modes remains weak. For a resonant monocromatic drive, this typically holds for a drive strength smaller than the driven mode decay rate $\Omega\ll\Gamma_{\alpha}$.
@@
Additionally, in the subwavelength regime in which the emitters spacing is smaller than their transition wavelength, modes with decay rates much smaller than independent emitters $\Gamma_{\alpha}\ll\Gamma$ appear, which are called subradiant [25, 4]. These arise due to destructive interference of the field radiated by different emitters. Despite the long lifetimes, by the same interference mechanism, subradiant modes are also linearly decoupled from far-field drives, thus they are not excited under a weak driving field [4, 27, 67]. We remark that this discussion is not specific to the case of arrays, but also extends to disordered ensembles, often realised in cold-atoms experiments [14, 15, 19, 23, 26] – also described by the model (1)-(3) upon specifying the appropriate couplings $V_{ij}$ and $\Gamma_{ij}$.
@@
For these reasons, even in the subwavelength regime and in the presence of – in principle, highly susceptible – subradiant states, the weak-drive regime has been widely regarded as a linear regime that is accurately described by classical equations of motion [53, 29, 36, 66, 8, 9, 45, 63, 49, 24, 26], equivalent to the noninteracting bosons description. These were expected to become exact descriptions in the limit of vanishing drive intensity [53, 29, 36, 66], without any distinction being made between the cases of disordered and ordered ensembles. This conclusion turns out to be wrong in the case of arrays of large sizes.
@@
In this case, discrete translational invariance implies that lattice momentum k is a good quantum number, and the bosonic eigenmodes correspond to reciprocal-space modes $a_{\textbf{k}}=\sum_{j}{\rm e}^{{\rm i}\textbf{k}\cdot\textbf{ r}_{j}}a_{j}/\sqrt{N}$, with dispersion relation and decay rates given by the Fourier transforms of (4) and (5) $V_{\textbf{k}}$ and $\Gamma_{\textbf{k}}$. These are shown in Fig. 1 (b) for our geometry (for $N\rightarrow\infty$). In the subwavelength regime, an entire manifold of subradiant eigenmodes arises, with decay rates strongly suppressed by the number of emitters. These decay rates have been found [4] to depend polynomially on $N$ for open boundary systems $\Gamma_{\textbf{k}}\sim N^{-\alpha}$, with $\alpha$ seemingly universal for the most subradiant modes with $\alpha=3$ in 1D, and $\alpha=6$ for a 2D square lattice. It was also shown that the residual decay is mainly due to scattering off the array boundaries, and that for periodic boundary conditions in 1D the suppression becomes exponential $\Gamma_{\textbf{k}}\sim\exp(-N)$.
@@
This suppression can be understood by subradiant modes having an energy much smaller than free-space photons with the same longitudinal momentum – the momentum along the array –, thus their excitations not to be able to decay into single photons conserving energy and momentum [4, 27]. From this argument, it follows that this happens whenever the mode dispersion intersects the light-cone within the Brillouin zone [4], corresponding to the condition $k_{0}a\leq\pi$ in 1D and $k_{0}a\leq\sqrt{2}\pi$ for a 2D square lattice, and that those modes appear at large momenta $\left\lvert\textbf{k}\right\rvert>k_{0}$, as in the case shown in Fig. 1 (b).
@@
Perturbative considerations. While the direct coupling of a far-field drive with subradiant modes is suppressed, resonant and efficient excitation of these modes can still happen through higher-order processes in perturbation theory, mediated by the nonlinearities of the atoms.
@@
These processes can be identified by expanding a Holstein-Primakoff representation of the spin in terms of bosons (see SM Note F), under the assumption that the drive is sufficiently weak, such that the steady state has a low density of excitations around the state with all spins pointing down. One can then expand simultaneously in the drive and in the nonlinear terms, and retain only resonant (i.e. energy-conserving) processes – as non-resonant ones are energetically suppressed. For our geometry – with a drive direction orthogonal to the array –, the lowest-order such processes involve the pair of subradiant modes with momenta $k$ and k-k, such that $V_{\pm k}=\Delta$. These read $(\Omega/\sqrt{N}a^{\dagger}_{-k}a^{\dagger}_{k}a_{k=0})_{t_{2}}({\sqrt{N}\Omega}a^{\dagger}_{k=0})_{t_{1}}$ and $({V_{k}}{/N}a^{\dagger}_{-k}a^{\dagger}_{k}a_{k=0}a_{k=0})_{t_{3}}({\sqrt{N}\Omega}a^{\dagger}_{k=0})_{t_{2}}({\sqrt{N}\Omega}a^{\dagger}_{k=0})_{t_{1}}$, where $a$ is the boson annihilation operator. These processes involve an intermediate (virtual) excitation of the $k=0$ mode, and convert two drive photons (as by the $\Omega^{2}$ factors) into a pair of subradiant excitations, conserving energy and momentum (see the illustration in panel (b) of Fig. 1). Importantly, lattice-momentum conservation ensures that the matrix element of these processes (between e.g. an initial vacuum state and the subradiant eigenstates) can also be non-negligible – a consideration that is distinct from energetics. We note that this would not be the case in the case of disordered ensembles, to which we therefore expect our conclusions not to apply – on the contrary we expect a similar phenomenology in arrays of other geometries.
@@
To estimate the effect of the above processes, we first treat them perturbatively with respect to the noninteracting theory – the single-particle bosonic model with an added coherent drive, equivalent to the classical approaches of [53, 29, 36, 66, 8, 9, 45, 63, 49, 24, 26] –, assuming a weak driving $\Omega$ and a large number of emitters $N$ (see SM Note A). In this respect, the fluctuations of the $k=0$ mode — which is directly, although off-resonantly driven — around its average value can be safely neglected. Keeping only the resonant pair of subradiant modes (on the same basis as earlier), this leads to an effective model for them. The noninteracting part of such model has Hamiltonian $H_{k}=\delta_{k}(a_{k}^{\dagger}a_{k}+a_{-k}^{\dagger}a_{-k})+{\lambda}_{k}(a_{k}^{\dagger}a_{-k}^{\dagger}+\text{h.c.})$, where coefficients $\delta_{k},\lambda_{k}$ are independent of $N$ and of leading order $\Omega^{2}$ in the drive strength. The dissipator, at leading order in $\Omega$ and $N$, is simply obtained by replacing $\sigma^{-}\rightarrow a$ in (3) and is diagonal in momentum, where further corrections are negligible since the decay rates of subradiant modes are already suppressed for large $N$.
The noninteracting part of the effective model obtained is then a model of a parametric amplifier [58]. In addition, in our problem the modes involved – owing to their subradiant nature – have decay rates $\Gamma_{\pm k}$ that decrease to zero with $N$. It is a well-known fact that a parametric amplifier develops a parametric instability for $\Omega\gtrsim\Gamma_{\pm k}^{1/2}$ [58], entailing a divergence of its steady-state moments. Then, in the case of our problem and for the most subradiant modes, this instability takes place for $\Omega\gtrsim N^{-\alpha/2}$ for open boundary conditions, and $\Omega\gtrsim\exp(-N/2)$ for periodic ones. Beyond such thresholds, the noninteracting model is dynamically unstable, and it therefore cannot be used as the basis of a perturbative analysis. This reveals that the steady-state of the full interacting problem (1) is non-perturbative, and that the widely-used noninteracting theories [53, 29, 36, 66, 8, 9, 45, 63, 49, 24, 26] are inadequate to describe large arrays of quantum emitters – even at a qualitative level, except at extremely-weak driving strengths. In particular, in the thermodynamics limit of $N\rightarrow\infty$, the threshold vanishes $\Omega\rightarrow 0$, and there is no linear regime left at weak driving strength. These conclusions motivate, and are supported by, the non-perturbative analysis discussed in the following.
@@
A non-perturbative analysis. To capture the steady state in the nonlinear (many-body) weak-drive regime discussed, we use the non-perturbative Dynamical Mean-Field Theory (DMFT) for Markovian spin systems introduced in [55, 54], for homogeneous steady states and in the thermodynamic limit $N\rightarrow\infty$. This is a bosonic version of DMFT, that maps the lattice model (1) onto an effective model of a single lattice site, coupled to an effective classical field and to a non-Markovian environment (see SM Note B), which are self-consistently determined, where the former is the analogous field of a Gutzwiller mean-field theory and the latter captures correlations beyond it. DMFT becomes exact in the noninteracting limit (as well as in the limit of independent emitters; see SM Note D), which is most relevant in the weak-drive regime (at the same time, we recall that here the $\Omega\rightarrow 0$ limit is non-perturbative, since we operate at $N\rightarrow\infty$).